OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (-6, 3).

OpenStudy (skullpatrol):

Hi again :-)

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (skullpatrol):

that depends on hard YOU want to try...

OpenStudy (anonymous):

ok.

OpenStudy (pawanyadav):

In this question You should know the property that product of slope of two perpendicular line is=-1.

OpenStudy (anonymous):

yeah

OpenStudy (pawanyadav):

Did this help you?

OpenStudy (anonymous):

not really

OpenStudy (skullpatrol):

did you make a sketch?

OpenStudy (pawanyadav):

Tell me according to the property, What should be the slope of required line.

OpenStudy (anonymous):



OpenStudy (anonymous):

Slope is \[\frac{ 3 }{ 2 }\]

OpenStudy (skullpatrol):

why ( 4, -2 )? the question gives ( -6, 3 )

OpenStudy (anonymous):

oh thats right i was looking at another question

OpenStudy (anonymous):

ugh

OpenStudy (skullpatrol):

try again :)

OpenStudy (pawanyadav):

Slope of required line is -2/3

OpenStudy (pawanyadav):

m1×m2=-1 m1=3/2 So. m2=-1/m1=-2/3

OpenStudy (pawanyadav):

Once. You get the slope the remaining part is similar to the earlier question.